Mike’s Cuisine Inc is a medium sized local restaurant specializing in South American Cuisine. The restaurant is situated close to a large state university. The owner, Mike, opened the business seven years ago, targeting the growing student population. However, the restaurant’s rate of growth has begun to decline recently, owing to increased competition. Mike’s reaction to the slow decline is to try and reduce demand volatility and improve his service package. Currently the restaurant is almost empty during early hours of operation, but later in the day, the waiting time for tables becomes excessive. Mike is considering entering the catering business. He is concerned that the catering business is highly cyclical and inherently riskier than operating a restaurant. He just paid $13,000 for market research, which projected revenue of $450,000 during the first year of operations. Launching the catering business, however, requires an investment of $320,000 for vehicles and equipment, plus another $75,000 for kitchen modifications. These assets would be depreciated on a straight-line basis to zero over a five-year period. Mike owns commercial property adjacent to his restaurant that according to an appraisal, could be sold for $680,000, after commissions and taxes. He must forego the sale of this property in order to enter the catering business. Mike worries about the fixed interest expense associated with partly financing his catering business through debt. He feels that the systematic risk of investinghis own money in that business is the same as investing in a security with a beta of 3.2. Mike’s Cusine has experienced the following average costs as a percentage of sales: 25% for cost of goods sold; 35% for variable costs, and 10% for net working capital. Mike anticipates the same cost profile for the catering business with the exception of a one-time, upfront investment of $50,000 in net working capital. He intends to finance 70% of the total investment with his own money and 30% with a bank loan at an effective annual interest rate of 8%. His businesses are taxed at a marginal tax rate of 35%. Mike plans to use his catering truck during football season to sell boxed meals that will contain a touchdown tailgate coupon for discounts on post-game dinners at Mike’s Cuisine. Mike is delighted by an announcement that the nearby large university is inviting bids on a catering contract for university events.

Questions:

1. If Mike enters the catering business, calculate the amount for his initial investment

2. Calculate the amount of the tax shield from depreciation during the first year of operations?

3. In deciding whether or not to enter the catering business, how relevant is the $13,000 spent on the marketing research?

4. Mike is concerned that he might need to sell more meals than expected per year to break even. Under which of the following conditions would Mike’s break even point increase?

a.Mike’s cost of goods sold decreases to 20%

b.Mike’s marginal tax rate increases by 5%

c.Mike’s of cost of goods sole increases to 40%

d.Mike’s investment in networking capital decreases by 10%

5. The Touchdown tailegate coupon promotion is known as a:

a. Consumer focused sales promotion

b.Consumer focused display advertisement

c.Trade-focused display advertisement

d.Trade focused sales promotion

6.Which of the following would most effectively help Mike capture lost sales during peak hours?

a.Providing entertainment to waiting customers

b.Standardizing the queuing procedures

c.Expanding the size of the facility

d. Shifting demand to a slower period

7.Which of the following would most elements of Mike’s Cuisine service package will most influence customer loyalty?

a.Explicit service

b.Implicit service

c.Facilitating goods

d.Location of service

8.Compared to Mike’s traditional consumers, the university is more likely to

a.Rely on emotional than objective criteria in selecting a caterer

b.Purchase in larger volume and expect lower negotiable prices

c.Purchase in larger volumes and expect higher non-negotiable prices

d.Rely on a single person to make the purchase decision
Answer #1

1). Initial investment = investment on vehicles and equipment + investment on kitchen modifications + opportunity cost of not selling the commercial property + one-time upfront investment in net working capital

= 320,000 + 75,000 + 680,000 + 50,000 = 1,125,000

2). Depreciation in year 1 = (investment on vehicles and equipment + investment on kitchen modifications)/5 =

(320,000 + 75,000)/5 = 395,000/5 = 79,000

Tax shield = depreciation*Tax rate = 79,000*35% = 27,650

3). $13,000 spent on market research is a sunk cost and will not be taken into account for deciding whether to enter the catering business or not.

4). Option c is correct.

break -even point Q = (Price per unit - variable cost per unit)/fixed cost

So, Q will increase when cost of goods sold increases.

5). Option a is correct. The Touchdown tailgate coupon promotion is a customer focused sales promotion as it is aimed at attracting customers by offering coupons.

6). Option b is correct. Standardizing queuing procedures will help in cutting down on excessive wait time. However, that can work only up to a point. If after standardizing, wait time is still too much then expansion of the facility will be needed.

7). Option a is correct. Customer loyalty will be influenced mostly by explicit service such as quality of the meal, attitude of the servers, etc.

8). Option b is correct. For catering contracts, hiring parties prefer to purchase in larger volume and expect lower negotiable prices.